a few questions:
1.Give a neighbourhood of a point 7.3
2. Is [0,1] an open set?
1
Expert's answer
2020-04-06T17:55:51-0400

We will use the following definition of a base of topology: every nonempty open set is an union of elements from the base.

If we consider the standard topology on "\mathbb R" (that is topology, base "\mathcal B" of which is the set of open intervals "\{(a,b)|a<b\}", then

1)"\bigcup\limits_{z\in\mathbb Z} \left(z,z+\frac{1}{2}\right)" is a neighbourhood of "7.3", because "7.3\in(7,7.5)\subset\bigcup\limits_{z\in\mathbb Z} \left(z,z+\frac{1}{2}\right)" and "\bigcup\limits_{z\in\mathbb Z} \left(z,z+\frac{1}{2}\right)" union of open intervals.

"(7,7.5)" is also a neighbourhood of "7.3", because "7.3\in(7,7.5)" and "(7,7.5)" is an open interval.

2)Prove that if a base "\mathcal B" of topology is given, then equivalent definition of open sets is the following:

Nonempy "A" is open set if and only if for every "x\in A" there is "U_x\in\mathcal B" such that "x\in U_x" and "U_x\subset A".

Let for every "x\in A" there is "U_x\in\mathcal B" such that "x\in U_x" ​ and "U_x\subset A" . Then we have "A=\bigcup\limits_{x\in A}\{x\}\subset\bigcup\limits_{x\in A}U_x" and, since "U_x\subset A" for every "x\in A", we have "\bigcup\limits_{x\in A}U_x\subset A". So "A=\bigcup\limits_{x\in A}U_x", that is "A" is open set by definition of a base of topology.

Now let "A" be an nonempty open set. Take arbitrary "x\in A". By the definition of a base of topology we have "A=\bigcup\limits_{Y\in\mathcal K}Y" for some "\mathcal K\subset\mathcal B". Since "x\in A", there is "Y_0\in K" such that "x\in Y_0". Since "\mathcal K\subset\mathcal B", we have "Y_0\in\mathcal B". So "x\in Y_0", where "Y_0\in B" and "Y_0\subset A".


Hence for the standard topology on "\mathbb R" we have "A" is an nonempty open set if and only if for every "x\in A" there is an open inerval "(c,d)" such that "x\in (c,d)" and "(c,d)\subset A".

Prove that "[0,1]" is not an open set. Indeed, "0\in[0,1]", but for every open interval "(c,d)", which contains , we have "\frac{c}{2}\in(c,d)\setminus [0,1]", that is "(c,d)\not\subset [0,1]". So "[0,1]" is not an open set.


Answer: 1)"\bigcup\limits_{z\in\mathbb Z} \left(z,z+\frac{1}{2}\right)", "(7,7.5)" are neighbourhoods of "7.3" in stanard topology on "\mathbb R"

2)"[0,1]" is not an open set in stanard topology on "\mathbb R".